Can i need ur help for this problem....thank you for the help.....😊

4x+7y=-2
6x-8y=5

Using linear equition substitution and illimination

Answer by rothauserc(4718)

1)4x+7y = -2
2) 6x-8y = 5
:
solve equation 2 for x
:
6x = 8y+5
x = (8y+5)/6
:
substitute for x in equation 1
:
4[(8y+5)/6] +7y = -2
:
(32y+20)/6 +7y = -2
:
multiply both sides of = by 6
:
32y+20 +42y = -12
:
74y = -32
:
*********************************
y = -32/74 = -16/37 approx −0.4324
x = [8(-16/37)+5]/6 approx 0.2568
:
check answer
:
1) 4*(0.2568) +7*(-0.4324) = -2
-1.9999 approx -2
:
2) 6*(0.2568) -8*(-0.4324) = 5
5.0002 approx 5
:
our answer checks
**********************************
: